My problem states: A nitrogen gas laser pulse with a wavelength of 337 nm contains 3.83 mJ of energy. How many photons does it contain? When I plug in the equations to my TI89 Titanium calculator, the answer I get is 6.49x10^-23 but your answer and the answer in my text is 6.49x10^15. Why? What am I plugging in wrong?
Answer #1

Wavelength =337nm =337×10-9m

Energy E=hc/wavelength

E=3.83mJ=3.83×10-3J

Energy of single photon is

E=hc/λ

E=(6.626×10-34 J s×3×108 m/s)/(337×10-9m)

=5.90×10-19J

Number of photons in the laser is

n=Total Energy/Energy per photon

n=(3.83×10-3J)/(5.90×10-19J/photon) =6.49×1015photons

You took 5.90×10-19 as 5.90×1019 in the calculation of n

So you got 6.49×10-23 as answer

But answer is 6.49×1015.

Thank you.

​​​